OpenStudy (anonymous):

what is the slope of 1,2.5 and -1,-2.5

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

-2.5-2.5/-1-1=5/2

OpenStudy (anonymous):

\[\frac{\frac{25}{10}-\left(-\frac{25}{10}\right)}{1-(-1)}=\frac{5}{2} \]

OpenStudy (whpalmer4):

The slope is the difference in the y values divided by the difference in the x values (you have to keep them in the same order!): \[m = \frac{y_2-y_1}{x_2-x_1}\] In your case, \[m=\frac{2.5-(-2.5)}{1-(-1)} = \frac{5}{2}\]

OpenStudy (whpalmer4):

It also works to do them in the other order: \[m=\frac{-2.5-2.5}{-1-1}=\frac{-5}{-2} = \frac{5}{2}\]

OpenStudy (whpalmer4):

You just can't use the x values in one order, and the y values in the other: \[m=\frac{-2.5-2.5}{1-(-1)}=\frac{-5}{2}\]which is incorrect.